OpenStudy (anonymous):

find the x intercep of the graph of the equation y=-3x-1

OpenStudy (amistre64):

y=0 when x=?

OpenStudy (amistre64):

0 = -3x-1 1 = -3x x = -1/3

OpenStudy (anonymous):

so x = -1/3 and y =0

OpenStudy (amistre64):

yes